OpenStudy (anonymous):

Simplify completely quantity 8 x minus 16 over quantity x squared minus 13 x plus 22 and find the restrictions on the variable.

OpenStudy (anonymous):

quantity x minus 2 over quantity x minus 11, x ≠ 11 quantity x minus 2 over quantity x minus 11, x ≠ 2, x ≠ 11 8 over quantity x minus 11, x ≠ 11 8 over quantity x minus 11, x ≠ 2, x ≠ 11

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

8 x minus 16 = 8(x -2) x squared minus 13 x plus 22 x^2 -13x + 22 what are the factors of +22 such that when you add you get -13 use -11 and -2 (x-11)(x-2) 8 x -16 / x^2 -13 x + 22 = 8(x-2)/ (x-11)(x-2) cancel out (x-2) = 8/(x-11) option D